# Scoreformer: A Surrogate Model For Large-Scale Prediction of Docking Scores

### v1

@@ Abstract
Álvaro Ciudad1, *, Adrián Morales-Pastor1, Laura Malo2, 3, Isaac Filella-Mercè4, Victor Guallar3, 4, Alexis Molina1, 3, *
1 Department of Artificial Intelligence, Nostrum Biodiscovery S.L., Barcelona, Spain
2 Department of Drug Discovery, Nostrum Biodiscovery S.L., Barcelona, Spain
3 Electronic and Atomic Protein Modelling Group, Barcelona Supercomputing Center, Barcelona, Spain
4 Institució Catalana de Recerca i Estudis Avançats (ICREA), Barcelona, Spain
* To whom correspondence may be addressed: {alvaro.ciudad, alexis.molina}@nostrumbiodiscovery.com
@@ 1 Introduction
Virtual screening (VS) is a key computational technique in drug discovery, utilized for evaluating the potential binding affinity of numerous molecules with target proteins. The main aim of VS is to identify molecules with potential interaction capabilities, thereby streamlining the drug discovery process and reducing the need for extensive, costly experimental assays (Lavecchia and Di Giovanni, 2013). High throughput virtual screening (HTVS) represents an advanced and more efficient form of VS. Characterized by its ability to rapidly assess vast molecular libraries, HTVS significantly contributes to identifying potential biological interactions.
The evolution of HTVS has been significantly influenced by the advancement of combinatorial chemistry, enabling the generation of extensive and diverse compound libraries. Previously, drug discovery relied on smaller and less diverse libraries, limiting the scope of potential candidates. The integration of combinatorial chemistry with HTVS allows for the utilization of larger compound libraries, alongside advanced computational evaluation techniques. This integration has accelerated the process of identifying suitable drug candidates by enabling rapid and precise assessment of ligand affinity across these vast libraries. Such efficiency is crucial in modern drug discovery as it can considerably shorten the hit-finding phase and enhance the overall hit rate. As a result, there’s an increasing need for methodologies that can swiftly and reliably navigate this expanded chemical space (Sadybekov and Katritch, 2023).
@@ 3.1 ScoreFormer Architecture
$\mathbf{X}_{i}^{(t+1)}=\text{U}^{(t)}\left(\mathbf{X}_{i}^{(t)},\underset{(j,i)\in E}{\bigoplus}\left(\text{M}^{(t)}\left(\mathbf{X}_{i}^{(t)},E_{j\rightarrow i},\mathbf{X}_{j}^{(t)}\right)\right)\right)$ (1)
@@ 4.1 Prediction of docking scores
In implementing a graph transformer model like ScoreFormer, one of our primary objectives was to achieve superior generalization in predicting docking scores. This capability can be relevant in drug discovery campaigns, where there is a significant interest in compounds across different ranges of docking scores. Such comprehensive predictive ability is beneficial not only for HTVS tasks but also for parallel applications like Quantitative Structure-Activity Relationships (QSAR) (Hansch and Fujita, 1964).
@@ 4.2 Molecular hit recovery
ScoreFormer’s ability to recover molecular hits, crucial for its application in prospective HTVS campaigns, was thoroughly evaluated using the metrics stated in Appendix C. The analysis, as illustrated in Table 1, demonstrated ScoreFormer’s superior performance in accurately identifying and prioritizing compounds with high binding affinity across various datasets and conditions. This proficiency is especially evident when comparing its performance to FiLMv2, a benchmark model in the field. ScoreFormer consistently outperformed FiLMv2 across multiple recovery metrics. Due to its increase in the number of parameters, the model is able to capture the topological features responsible for good docking scores along the different systems and constrain set. These results highlight ScoreFormer’s robustness and reliability in hit recovery. Notably, L-ScoreFormer also displayed commendable performance, often closely following the FiLMv2 architecture. This observation underscores the effectiveness of both of our architectures in molecular hit recovery tasks. The consistent results of both ScoreFormer and L-ScoreFormer across diverse screening scenarios reinforce the versatility of our approach. See further results at Appendix E.
@@ 4.3 Inference speed
Model Samples/s 128M time (h) Speedup # parameters
FiLMv2 1323.942 26.850 1.000 102977
ScoreFormer 2186.828 16.259 1.652 5398273
L-ScoreFormer 2468.404 14.404 1.864 147457
@@ A.1 System preparation
All systems were prepared with the same protocol and the structures were the following: CDK2 with PDB id 3BHV [Bettayeb et al., 2007], D4 with PDB id 5WIU [Wang et al., 2017] and HIV-1 with PDB id 3AID [Rutenber et al., 1996].
Protonation states were assigned at pH 7.4 using the Protein Preparation Wizard [Madhavi Sastry et al., 2013]. Hydrogen bonds were optimized with PROPKA [Søndergaard et al., 2011] and a restrained minimization using the atom force field OPLS4 [Lu et al., 2021] was applied to each system.
@@ A.2 Glide docking
    CDK2: hydrogen bond constraint to the central donor of the CDK2 hinge region (LEU83) [Li et al., 2015].
@@
    D4: hydrogen bond constraint to bias the selection of the compounds to selective agonists (ASP115) [Wang et al., 2017].
@@
    HIV-1: hydrogen bond constraint to the highly conserved aspartate in an aspartic protease (ASP25) active site to mimic ligand interactions of known inhibitors [Rutenber et al., 1996].
@@ C.3 Area under recall threshold curve ($AURTC_{\zeta}$)
The recall threshold curve is defined as a slice of the RES by fixing the threshold $\zeta$ (the ratio of the model predictions to be considered positive). As in Hosseini et al. [2022], we report AURTC using two $\zeta$ values for all experiments: 0.01 and 0.001.

### v2

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
Virtual screening (VS) is a key computational technique in drug discovery, utilized for evaluating the potential binding affinity of numerous molecules with target proteins. The main aim of VS is to identify molecules with potential interaction capabilities, thereby streamlining the drug discovery process and reducing the need for extensive, costly experimental assays (Lavecchia & Di Giovanni, 2013). HTVS represents an advanced and more efficient form of VS. Characterized by its ability to rapidly assess vast molecular libraries, HTVS significantly contributes to identifying potential biological interactions.
The evolution of HTVS has been significantly influenced by the advancement of combinatorial chemistry, enabling the generation of extensive and diverse compound libraries. Previously, drug discovery relied on smaller and less diverse libraries, limiting the scope of potential candidates. The integration of combinatorial chemistry with HTVS allows for the utilization of larger compound libraries, alongside advanced computational evaluation techniques. This integration has accelerated the process of identifying suitable drug candidates by enabling rapid and precise assessment of ligand affinity across these vast libraries. Such efficiency is crucial in modern drug discovery as it can considerably shorten the hit-finding phase and enhance the overall hit rate. As a result, there’s an increasing need for methodologies that can swiftly and reliably navigate this expanded chemical space (Sadybekov & Katritch, 2023).
@@ 3.1 ScoreFormer Architecture
$\mathbf{X}_{i}^{(t+1)}=\text{U}^{(t)}\left(\mathbf{X}_{i}^{(t)},\underset{(j,i)\in E}{\bigoplus}\left(\text{M}^{(t)}\left(\mathbf{X}_{i}^{(t)},E_{j\rightarrow i},\mathbf{X}_{j}^{(t)}\right)\right)\right)$
@@ 4.1 Prediction of docking scores
In implementing a graph transformer model like ScoreFormer, one of our primary objectives was to achieve superior generalization in predicting docking scores. This capability can be relevant in drug discovery campaigns, where there is a significant interest in compounds across different ranges of docking scores. Such comprehensive predictive ability is beneficial not only for HTVS tasks but also for parallel applications like Quantitative Structure-Activity Relationships (QSAR) (Hansch & Fujita, 1964).
@@ 4.2 Molecular hit recovery
ScoreFormer’s ability to recover molecular hits, crucial for its application in prospective HTVS campaigns, was thoroughly evaluated using the metrics stated in Appendix C. The analysis, as illustrated in Table 1, demonstrated ScoreFormer’s superior performance in accurately identifying and prioritizing compounds with high binding affinity across various datasets and conditions. This proficiency is especially evident when comparing its performance to FiLMv2, a benchmark model in the field. ScoreFormer consistently outperformed FiLMv2 across multiple recovery metrics. Due to its increase in the number of parameters, the model is able to capture the topological features responsible for good docking scores along the different systems and constrain set. These results highlight ScoreFormer’s robustness and reliability in hit recovery. Notably, L-ScoreFormer, also displayed commendable performance, often closely following the FiLMv2 architecture. This observation underscores the effectiveness of both of our architectures in molecular hit recovery tasks. The consistent results of both ScoreFormer and L-ScoreFormer across diverse screening scenarios reinforce the versatility of our approach. See further results at Appendix E.
@@ 4.3 Inference speed
Model Samples/s 128M time (h)
FiLMv2 1323.942 26.850
ScoreFormer 2186.828 16.259
L-ScoreFormer 2468.404 14.404
Model Speedup # parameters
FiLMv2 1.000 102977
ScoreFormer 1.652 5398273
L-ScoreFormer 1.864 147457
@@ A.1 System preparation
All systems were prepared with the same protocol and the structures were the following: CDK2 with PDB id 3BHV (Bettayeb et al., 2007), D4 with PDB id 5WIU (Wang et al., 2017) and HIV-1 with PDB id 3AID (Rutenber et al., 1996).
Protonation states were assigned at pH 7.4 using the Protein Preparation Wizard (Madhavi Sastry et al., 2013). Hydrogen bonds were optimized with PROPKA (Søndergaard et al., 2011) and a restrained minimization using the atom force field OPLS4 (Lu et al., 2021) was applied to each system.
@@ A.2 Glide docking
    CDK2: hydrogen bond constraint to the central donor of the CDK2 hinge region (LEU83) (Li et al., 2015).
@@
    D4: hydrogen bond constraint to bias the selection of the compounds to selective agonists (ASP115) (Wang et al., 2017).
@@
    HIV-1: hydrogen bond constraint to the highly conserved aspartate in an aspartic protease (ASP25) active site to mimic ligand interactions of known inhibitors (Rutenber et al., 1996).
@@ C.3 Area under recall threshold curve ($AURTC_{\zeta}$)
The recall threshold curve is defined as a slice of the RES by fixing the threshold $\zeta$ (the ratio of the model predictions to be considered positive). As in Hosseini et al. (2022), we report AURTC using two $\zeta$ values for all experiments: 0.01 and 0.001.